OpenStudy (anonymous):

What are the classifications of each system? A. consistent independent B. coincident C. inconsistent 1. 4x-2y=10 2x-y=5 2. y=2x-3 -2x+y=-5 3. 2x-5y=14 3x+4y=10 4. y=4x-3 12x-3y=9

OpenStudy (amistre64):

it worked :)

OpenStudy (amistre64):

the first thing we need to do is to know what the definitions for ABC are in order to apply them

OpenStudy (anonymous):

I know what they are consistent independent- one solution coincident-infinite solutions inconsistent- no solutions

OpenStudy (amistre64):

now, this question assumes you know how to see the properties of a line from the equations if one equation can be multiplied to get the other equation, then they are the same line if one equation can be multiplied to get the other equation (EXCEPT for the constant), then they are parallel lines if neither of these hold, then they have to cross at 1 point

OpenStudy (anonymous):

This I learned today, but I just don't know what the equations are, like what the equation is in line form.

OpenStudy (amistre64):

lets take the first one, and test it out 4x - 2y = 10 2x - y = 5 in order to get the equations looking somewhat the same, then 2x times something has to equal 4x ... would you agree that multiplying the bottom by 2 would accomplish this?

OpenStudy (anonymous):

Yes, I would.

OpenStudy (anonymous):

So would it be A?

OpenStudy (amistre64):

4x - 2y = 10 2(2x - y = 5) 4x - 2y = 10 4x - 2y = 10 they are the same equation, therefore the same line coincident

OpenStudy (amistre64):

y=2x-3 -2x+y=-5 i would suggest getting all the x and y parts to the same side, what do we get from that?

OpenStudy (anonymous):

y=2x-5

OpenStudy (amistre64):

getting them both in y= form .. thats fine to :) y=2x-3 y=2x-5 so, are they the same? or do they differ only by the constant? or neither?

OpenStudy (anonymous):

3 and 5 so, they differ by the constant

OpenStudy (amistre64):

then they are the same line, but in different positions ... which option is it?

OpenStudy (anonymous):

I think C

OpenStudy (amistre64):

parallel lines, so yes .. they have no points in common

OpenStudy (amistre64):

2x-5y=14 3x+4y=10 try this one out ...

OpenStudy (anonymous):

Perhaps B

OpenStudy (amistre64):

show me your 'proof" try to turn one into the other ... or at best one variables coeff into the other

OpenStudy (anonymous):

Well, I don't think you can turn one into the other

OpenStudy (amistre64):

i dont think so either, but its fun to try :) 3(2x-5y=14)/2 3x+4y=10 3(x-5/2y=7) 3x+4y=10 3x-15/2y=21 3x+4y=10 you cannot change one into a multiple of the other so they are neither parallel or coincident they must have only 1 point in common

OpenStudy (amistre64):

y=4x-3 12x-3y=9 you know the drill by now i hope :) see if we can change it

OpenStudy (anonymous):

So a?

OpenStudy (anonymous):

I'm not sure, but i think 12x-3=9 is like this in this problem -12x+9

OpenStudy (amistre64):

1. same line, same constant, coincident lines 4x-2y=10 2x-y=5 2. same line, different constants, parallel lines y=2x-3 -2x+y=-5 3. neither parallel or consistent ... crossing lines 2x-5y=14 3x+4y=10

OpenStudy (amistre64):

y=4x-3 12x-3y=9 get all the x y parts to the same side -4x+ y= -3 12x -3y=9 multiply thru by -3 and compare

OpenStudy (anonymous):

I'm not sure how to do that

OpenStudy (amistre64):

you are not sure how to multiply by -3?

OpenStudy (anonymous):

no what to multiply -3 by

OpenStudy (amistre64):

-4x+ y= -3 12x -3y=9 ^ ^ we have a -3y down here and should prolly have a -3y up top as well

OpenStudy (amistre64):

multiply the top part by -3

OpenStudy (anonymous):

So it would be identical to the other equation

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so if they are the exact same does that make it B?

OpenStudy (amistre64):

A. consistent independent , crossing lines B. coincident , identical C. inconsistent , off by a constant yes, B :)

OpenStudy (anonymous):

okay thanks!:)

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

Bye, Have a Great Rest of Your Day!:)